OpenStudy (anonymous):

Which has the greatest value? Answer 1 to the twelfth power 2 to the third power 9 to the first power fifteen to the zero power

OpenStudy (bacon):

well 1 times 1 itself is gonna always be 1 and 2x2x2 is 8 and 9 to the first power is gonna be 9 and anything to the zero power is 1

OpenStudy (anonymous):

\[1^{12}=1\times1\times1\times1\times1\times1\times1\times1\times1\times1\times1\times1=1\]\[2^3=2\times2\times2=8\]\[9^1=9\]\[15^0=1\] now out of 1,8,9, and 1, which would be the greatest? :)

OpenStudy (anonymous):

9 to the first power

OpenStudy (bacon):

yes

OpenStudy (anonymous):

Great! :)

OpenStudy (anonymous):

you're welcome....

OpenStudy (bacon):

can i get a meddle?

OpenStudy (anonymous):

Thank you yummy. :)

OpenStudy (anonymous):

@Bacon you mean a Medal*? lol

OpenStudy (bacon):

yes lol